OpenStudy (anonymous):

First time. Help! Solve the following system of equations and show all work: y=x^2+3 y=x+5

OpenStudy (anonymous):

y=x^2+3 y=x+5 We can do this x^2 + 3 = x + 5 x^2 - x + 3 - 5 = 0 x^2 - x - 2 = 0 Now , solve this equation for x .

OpenStudy (anonymous):

so would it be 2x-x-2=0 x-2=0 x=2?

OpenStudy (anonymous):

(x-2)(x+1) = 0 X-2 = 0 or x + 1 = 0 so x is either -1 or 2

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

there are no options

OpenStudy (anonymous):

Are we suppose to find y also?

OpenStudy (anonymous):

i don't know. it's not a multiple choice question :/ it just say solve the system of equations

OpenStudy (phi):

x= -1 or x=2 use those numbers to find the corresponding y in y=x+5

OpenStudy (phi):

this problem has two solutions i.e. two (x,y) points that lie on both the line and the parabola.

OpenStudy (anonymous):

so it doesn't matter which x i use ?

OpenStudy (phi):

you use both, one at a time. you want to find two (x,y) pairs where x is -1 in the first pair (and you need to find the y) and x=2 in the other pair

OpenStudy (anonymous):

so it would be (-1,4) or (2,7)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

what do i do about the other y equation?

OpenStudy (phi):

you could use either equation to find y but it makes sense to use the easier one. you are done! btw, here is what you found... where the line meets the parabola

OpenStudy (anonymous):

okay. thank you both so much! how do i give you both medals for helping me out?